The market demand function is

Q = 10000 - 1000p

Each firm has a marginal cost of m=​$0.28. Firm​ 1, the​ leader, acts before Firm​ 2, the follower. Solve for the​ Stackelberg-Nash equilibrium​ quantities, prices, and profits. Compare your solution to the​ Cournot-Nash equilibrium. The​ Stackelberg-Nash equilibrium quantities are

q1 = ____ units and q2= ____ units.  ​(Enter your responses as whole​ numbers.)

The Stackelberg-Nash equilibrium price is:

p=$_____________

Profits for the firms are

profit1=$_______________

and profit2=$_______________

The Cournot-Nash equilibrium quantities are:

q1=______________units

and q2=______________units

The Cournot-Nash equilibrium price is:

p=$______________

Profits for the firms are

profit1=$_____________ and profit2=$_______________
Answer #1

8 = 10000 – 1080p p= 10- Mc = 6.28 81= lender = (a 2XL looo - (10-0.28) 2X1000 = 9.72 Soo = 9.728 500 leader a. = 4860 s a 9Іe EasyBour Manage Chege Study Profila С ТА-ТО — A-30 % в Р 336 Ла-5 - — 8,%-) Cournot Ремо 10 - 15 = 6,28 (Фоо 9 +2 = Тръ8)